# Bilinear spherical maximal function with fractal dilations

### v1

@@ 5. Estimate for local maximal operator
Observe that using Minkowski’s integral inequality, we have
$\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)(x)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\leq$ $2^{|\mathbf{n}|\frac{\beta}{2}}\int_{{\mathbb{R}}^{d}}\bigg(\int_{I}\left|U_{s,N}^{n_{2}}f_{2}(x-y)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\frac{dy}{\left(1+4\pi^{2}\left|2^{n_{2}}y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}.$ (5.3)
We take $L^{2}$-norm in the above and use Plancherel’s theorem to get that
$\left\|\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{2}}\leq$ $2^{|\mathbf{n}|\frac{\beta}{2}}\left\|\bigg(\int_{I}\big|U_{s,N}^{n_{2}}f_{2}\big|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{2}}\int_{{\mathbb{R}}^{d}}\frac{dy}{\left(1+4\pi^{2}\left|2^{n_{2}}y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}$ (5.4)
$\lesssim$ $2^{-n_{2}d}2^{-|\mathbf{n}|\frac{1-\beta}{2}}\|f_{2}\|_{L^{2}}.$
Further, using the estimate 5.3 along with Lemma 4.2 and Minkowski’s integral inequality, we get that
$\left\|\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{1}}\lesssim 2^{-n_{2}\frac{d}{2}}2^{-|\mathbf{n}|\frac{1-\beta}{2}}\|f_{2}\|_{L^{1}}.$ (5.5)
Next, since $\widehat{\sigma}(s\xi,s\eta):=m(s\xi,s\eta)e^{2\pi is|(\xi,\eta)|}$ with $|m(\xi,\eta)|\lesssim|(\xi,\eta)|^{-\frac{2d-1}{2}}$, by an application of Lemma 4.1, we obtain
$\left\|\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{\mathbf{n}}_{s,N}\left(F_{N}f_{1}\right)(\cdot,\eta)\right|d\eta\right\|_{L^{2}}\lesssim 2^{-\frac{2d-1}{2}|\mathbf{n}|}2^{n_{2}d}\|F_{N}f\|_{L^{2}}.$ (5.6)
Using Hölder’s inequality along with the estimates (5.4), (5.5) and (5.6), we obtain
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{1}}\lesssim$ $2^{-|\mathbf{n}|\frac{2d-\beta}{2}}\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\|F_{N}f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}$
$\lesssim$ $2^{-|\mathbf{n}|\frac{2d-1-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}$
and
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\mathcal{G}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{2/3}}$ $\lesssim 2^{-|\mathbf{n}|\frac{d-\beta}{2}}\bigg(\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\|F_{N}f_{1}\|_{L^{2}}^{\frac{2}{3}}\|f_{2}\|_{L^{1}}^{\frac{2}{3}}\bigg)^{\frac{3}{2}}$
@@
As described earlier, similar arguments as above, yield the following bounds for square function involving the derivative term $\frac{\partial}{\partial s}\mathcal{A}_{s}$.
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\tilde{\mathcal{G}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{1}}\lesssim 2^{-|\mathbf{n}|\frac{2d-3-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}},$
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\tilde{\mathcal{G}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{2/3}}\lesssim 2^{-|\mathbf{n}|\frac{d-4-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}.$
Substituting the estimates obtained above in (5.1), we have
$\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{1}}\lesssim$ $2^{\frac{|\mathbf{n}|}{2}}2^{-|\mathbf{n}|\frac{2d-1-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}$
$+\left(2^{-|\mathbf{n}|\frac{2d-1-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}\right)^{\frac{1}{2}}\left(2^{-|\mathbf{n}|\frac{2d-3-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}\right)^{\frac{1}{2}}$
$\lesssim$ $2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}},$
and
$\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{2/3}}\lesssim 2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}.$
@@
$\left\|\mathcal{M}^{\mathbf{n}}_{0}(f_{1},f_{2})\right\|_{L^{2/3}}$ $\lesssim 2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
This completes the proof of Lemma 5.2. ∎
@@
Next, we consider the partition ${\mathbb{R}}^{d}=\cup_{Q\in\mathfrak{C}}Q$, where $Q\in\mathfrak{C}$ are cubes with unit sidelengths and sides parallel to co-ordinate axes. For $a<0,$ let $aQ$ denote the cube which is concentric with $Q$ and have sidelength $a>0$. With this we can write
@@
Let $I\in\mathcal{E}_{0}$ be any interval. Without loss of generality, we may assume that $I=[t_{0}-2^{-|\mathbf{n}|-1},t_{0}+2^{-|\mathbf{n}|-1}]$. For any $t\in I$ and $|y|\leq 1$, we have
@@
where we have used the $L^{1}\times L^{1}\to L^{1}$-boundedness of the single bilinear spherical average $\mathcal{A}_{t}$ obtained in [IPS] in the third step and Cauchy-Schwarz inequality in the fourth step of the equation above. This completes the proof of Lemma 5.3. ∎
@@
Recall that in order to complete the proof of Theorem 2.1 we are required to prove desired $L^{p}$-estimates for the operator
@@
Note that since the notion of Minkowski dimension as described in Definition 1.1 is uniform in $k$. Since we are concerned with Hölder exponents $(p_{1},p_{2},p)$, the operator $\mathcal{M}^{\mathbf{n}}_{k}$ satisfies the same bounds as those of $\mathcal{M}^{\mathbf{n}}_{0}$ proved in Lemma 5.2 and Lemma 5.3. Moreover, the bounds are uniform in $k$. Next, we use these estimates for each fixed scale to prove the following results for the operator $\mathcal{M}^{\mathbf{n}}$.
@@
Let $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
$\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\|_{L^{1}}\lesssim 2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}.$
@@
Let $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
@@
    $\left\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\right\|_{L^{2/3,\infty}}\lesssim|\mathbf{n}|^{3}2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}},$
@@
    $\left\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\right\|_{L^{2/3,\infty}}\lesssim|\mathbf{n}|^{3}2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}$,
  3. (3)
@@
Note that if we assume Lemma 6.1 and Lemma 6.2, the proof of Theorem 2.1 follows by arguments similar to those given for the operator $\mathcal{M}_{E_{0}}$.
@@
$\lesssim\sum_{k\in{\mathbb{Z}}}2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|\mathbf{R}_{k+n_{1}}f_{1}\|_{L^{2}}\|\mathbf{R}_{k+n_{2}}f_{2}\|_{L^{2}}$
$\leq 2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\left(\sum_{k\in{\mathbb{Z}}}\|\mathbf{R}_{k+n_{1}}f_{1}\|_{L^{2}}^{2}\right)^{\frac{1}{2}}\left(\sum_{k\in{\mathbb{Z}}}\|\mathbf{R}_{k+n_{2}}f_{2}\|_{L^{2}}^{2}\right)^{\frac{1}{2}}$
$\lesssim 2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}.$
@@
6.2. Proof of parts $(1)$ and $(2)$ of Lemma 6.2
Note that due to symmetry it is enough to prove part $(1)$. The proof is based on a suitable adaptation of the bilinear Calderón-Zygmund theory and delicate analysis involving the exceptional set which arises in the Calderón-Zygmund decomposition of input function. Without loss of generality we may assume that $\|f_{1}\|_{L^{1}}=\|f_{2}\|_{L^{2}}=1$. Given $\alpha>0$, we shall prove the following
$\Big|\{x:\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})(x)>\alpha\}\Big|\leq C|\mathbf{n}|^{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{-\frac{2}{3}}.$ (6.1)
We apply the Calderón-Zygmund decomposition to $f_{1}$ at height $2^{|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{\frac{2}{3}}$ to obtain the decomposition $f_{1}=g_{1}+h_{1}$ with a collection $\{Q_{\gamma}\}$ of disjoint cubes such that
$\|g_{1}\|_{L^{\infty}}\leq 2^{|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{\frac{2}{3}},~~h_{1}=\sum_{\gamma}h_{\gamma},~~supp(h_{\gamma})\subset Q_{\gamma},$
$\|h_{\gamma}\|_{L^{1}}\lesssim 2^{|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{\frac{2}{3}}|Q_{\gamma}|~~\text{with}~~\int_{Q_{\gamma}}h_{\gamma}=0,~~\text{and}$
$\sum_{\gamma}\big|Q_{\gamma}\big|\leq 2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{-\frac{2}{3}}.$
@@
$|\{x\in\mathbb{R}^{d}:\mathcal{M}^{\mathbf{n}}(g_{1},f_{2})(x)>\frac{\alpha}{2}\}|\lesssim\frac{2^{-|\mathbf{n}|\frac{2d-2-\beta}{2}}\|g_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}}{\alpha}.$
Since $\|g_{1}\|_{2}\lesssim 2^{|\mathbf{n}|\frac{d-3-\beta}{6}}\alpha^{\frac{1}{3}}$, we obtain
$|\{x\in\mathbb{R}^{d}:\mathcal{M}^{\mathbf{n}}(g_{1},f_{2})(x)>\frac{\alpha}{2}\}|$ $\lesssim 2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{-\frac{2}{3}}.$
@@
Then, $|\mathcal{F}|\leq 2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\alpha^{-\frac{2}{3}}$. Thus, by Chebyshev’s inequality, it is enough to show that
$\int_{\mathbb{R}^{d}\setminus\mathcal{F}}|\mathcal{M}^{\mathbf{n}}(h_{1},f_{2})(x)|^{\frac{2}{3}}~dx\lesssim|\mathbf{n}|^{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}.$ (6.2)
@@
$\int\limits_{\mathbb{R}^{d}\setminus\mathcal{F}}\Big|\mathcal{M}^{\mathbf{n}}_{k}(h^{i}_{1},f_{2})(x)\Big|^{\frac{2}{3}}dx\lesssim n_{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\min\big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\big\}\|h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|f_{2}\|^{\frac{2}{3}}_{L^{2}}.$
@@
$\int\limits_{\mathbb{R}^{d}\setminus\mathcal{F}}\Big|\mathcal{M}^{\mathbf{n}}_{k}(h^{i}_{1},f_{2})(x)\Big|^{\frac{2}{3}}dx\lesssim n_{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\min\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\}\|h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{\frac{2}{3}}_{L^{2}}.$
@@
$\lesssim\sum_{k,i\in{\mathbb{Z}}}n_{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\min\Big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\Big\}\|h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{\frac{2}{3}}_{L^{2}}$
$\lesssim n_{2}2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\sum_{k\in{\mathbb{Z}}}\bigg(\sum_{i\in{\mathbb{Z}}}\min\Big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\Big\}\bigg)^{\frac{1}{3}}$

### v2

@@ 5. Estimate for local maximal operator $\mathcal{M}_{E_{0}}$
Using Minkowski’s integral inequality along with Lemma 4.2, we get that
$\left\|\bigg(\sum_{I\in\mathcal{E}_{0}}\int_{I}\left|L_{n_{2}}\left(U_{s,N}^{n_{2}}f_{2}\right)\right|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{1}}\leq$ $2^{|\mathbf{n}|\frac{\beta}{2}}\left\|\bigg(\int_{I}\big|U_{s,N}^{n_{2}}f_{2}\big|^{2}\;ds\bigg)^{\frac{1}{2}}\right\|_{L^{1}}\int_{{\mathbb{R}}^{d}}\frac{dy}{\left(1+4\pi^{2}\left|2^{n_{2}}y\right|^{2}\right)^{\lceil\frac{d}{2}\rceil+1}}$ (5.11)
$\lesssim$ $2^{-n_{2}d}2^{-|\mathbf{n}|\frac{1-\beta}{2}}\|f_{2}\|_{L^{1}}.$
Next, by Minkowski’s integral inequality and an application of Lemma 4.1, we obtain
$\left\|\int_{\mathbb{R}^{d}}\sup_{1\leq s\leq 2}\left|B^{\mathbf{n}}_{s,N}\left(F_{N}f_{1}\right)(\cdot,\eta)\right|d\eta\right\|_{L^{2}}\lesssim 2^{-|\mathbf{n}|\frac{2d-1}{2}}2^{n_{2}d}\|F_{N}f\|_{L^{2}}.$ (5.12)
Using Hölder’s inequality along with the estimates (5.11) and (5.12), we obtain
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\mathcal{H}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{2/3}}$ $\lesssim 2^{-|\mathbf{n}|\frac{d-\beta}{2}}\bigg(\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\|F_{N}f_{1}\|_{L^{2}}^{\frac{2}{3}}\|f_{2}\|_{L^{1}}^{\frac{2}{3}}\bigg)^{\frac{3}{2}}$
@@
Similarly, we have the following bounds for the square function $\tilde{\mathcal{H}}^{\mathbf{n}}_{0,N}$.
$\Bigg\|\sum_{N=2^{n_{1}-1}}^{2^{n_{1}+1}}\tilde{\mathcal{H}}^{\mathbf{n}}_{0,N}(f_{1},f_{2})\Bigg\|_{L^{2/3}}\lesssim 2^{-|\mathbf{n}|\frac{d-4-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}.$
From (5.10), and $L^{p}$-estimates of $\mathcal{H}^{\mathbf{n}}_{0,N}$ and $\tilde{\mathcal{H}}^{\mathbf{n}}_{0,N}$ obtained above, we have
$\left\|\mathbb{I}^{\mathbf{n}}\right\|_{L^{2/3}}\lesssim 2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}.$ (5.13)
Combining the estimates in (5.9) and (5.13), we get
$\left\|\mathbb{I}^{\mathbf{n}}\right\|_{L^{2/3}}\lesssim\min\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{2}},2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}$
@@
$\left\|\mathbb{I}^{\mathbf{n}}\right\|_{L^{2/3}}$ $\lesssim\min\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{2}},2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\}\|f_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}}.$
Combining the above estimates for $\mathbb{I}^{\mathbf{n}}$ and similar estimates for $\mathbb{II}^{\mathbf{n}}$ along with (5.2) completes the proof of Lemma 5.2. ∎
@@
Next, we consider the partition ${\mathbb{R}}^{d}=\cup_{Q\in\mathfrak{C}}Q$, where $Q\in\mathfrak{C}$ are cubes with unit sidelength and sides parallel to co-ordinate axes. For $a>0,$ let $aQ$ denote the cube which is concentric with $Q$ and have sidelength $a$. With this we can write
@@
Let $I\in\mathcal{E}_{0}$ be any interval. Without loss of generality, we may assume that $I=[t_{0}-2^{-|\mathbf{n}|-1},t_{0}+2^{-|\mathbf{n}|-1}]$, where $t_{0}\in[1,2]$. For any $t\in I$ and $|y|\leq 1$, we have
@@
where we have used the $L^{1}\times L^{1}\to L^{1}$-boundedness of the single bilinear spherical average $\mathcal{A}_{t}$ obtained in [25] in the third step and Cauchy-Schwarz inequality in the fourth step of the equation above. This completes the proof of Lemma 5.3. ∎
@@
Recall that in order to complete the proof of Theorem 2.1 we are required to prove the desired $L^{p}$-estimates for the operator
@@
Note that by a change of variable $\xi\to 2^{k}\xi$ and $\eta\to 2^{k}\eta$, we have
$\mathcal{M}^{\mathbf{n}}_{k}(f_{1},f_{2})(x)=$ $\sup_{t\in E_{k}}\left|\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{d\sigma}(t(\xi,\eta))\widehat{\psi}\left(2^{-k-n_{1}}\xi\right)\widehat{\psi}\left(2^{-k-n_{2}}\eta\right)\widehat{f_{1}}(\xi)\widehat{f_{2}}(\eta)e^{ix\cdot(\xi+\eta)}~d\xi d\eta\right|$ (6.1)
$=$ $\sup_{s\in J}\left|\int_{{\mathbb{R}}^{d}\times{\mathbb{R}}^{d}}\widehat{d\sigma}(s(\xi,\eta))\widehat{\psi}\left(2^{-n_{1}}\xi\right)\widehat{\psi}\left(2^{-n_{2}}\eta\right)\widehat{(f_{1})_{2^{-k}}}(\xi)\widehat{(f_{2})_{2^{-k}}}(\eta)e^{i2^{k}x\cdot(\xi+\eta)}~d\xi d\eta\right|,$
where $s=2^{k}t\in 2^{k}E_{k}=J$ with $J\subset[1,2]$ and $(f)_{2^{-k}}=f(2^{-k}x)$.
Since the notion of Minkowski dimension as described in Definition 1.1 is uniform in $k$, the operator $\mathcal{M}^{\mathbf{n}}_{k}$ satisfies the same bounds as those of $\mathcal{M}^{\mathbf{n}}_{0}$ proved in Lemma 5.1, Lemma 5.2 and Lemma 5.3 for Hölder exponents $(p_{1},p_{2},p)$. Next, we use these estimates for each fixed scale to prove the following results for the operator $\mathcal{M}^{\mathbf{n}}$.
@@
Let $d\geq 1$ and $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
$\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\|_{L^{1}}\lesssim 2^{-|\mathbf{n}|\frac{2d-1-2\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}.$
@@
Let $d\geq 3$ and $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
@@
    $\left\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\right\|_{L^{2/3,\infty}}\lesssim|\mathbf{n}|^{3}\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{2}},2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\right\}\|f_{1}\|_{L^{1}}\|f_{2}\|_{L^{2}},$
@@
    $\left\|\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})\right\|_{L^{2/3,\infty}}\lesssim|\mathbf{n}|^{3}\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{2}},2^{-|\mathbf{n}|\frac{d-3-\beta}{2}}\right\}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{1}}$.
Lemma 6.3.
Let $d\geq 1$ and $0\leq\beta\leq 1$. Then for any $\mathbf{n}\in{\mathbb{N}}^{2}$, we have
@@
Note that if we assume Lemma 6.1. Lemma 6.2 and Lemma 6.3, the proof of Theorem 2.1 follows by arguments similar to those given for the operator $\mathcal{M}_{E_{0}}$.
@@
$\lesssim\sum_{k\in{\mathbb{Z}}}2^{-|\mathbf{n}|\frac{2d-1-2\beta}{2}}\|\mathbf{R}_{k+n_{1}}f_{1}\|_{L^{2}}\|\mathbf{R}_{k+n_{2}}f_{2}\|_{L^{2}}$
$\leq 2^{-|\mathbf{n}|\frac{2d-1-2\beta}{2}}\left(\sum_{k\in{\mathbb{Z}}}\|\mathbf{R}_{k+n_{1}}f_{1}\|_{L^{2}}^{2}\right)^{\frac{1}{2}}\left(\sum_{k\in{\mathbb{Z}}}\|\mathbf{R}_{k+n_{2}}f_{2}\|_{L^{2}}^{2}\right)^{\frac{1}{2}}$
$\lesssim 2^{-|\mathbf{n}|\frac{2d-1-2\beta}{2}}\|f_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}.$
@@
6.2. Proof of Lemma 6.2.
Note that due to symmetry, it is enough to prove part $(1)$. The proof is based on a suitable adaptation of the bilinear Calderón-Zygmund theory and delicate analysis involving the exceptional set which arises in the Calderón-Zygmund decomposition of input function. Without loss of generality we may assume that $\|f_{1}\|_{L^{1}}=\|f_{2}\|_{L^{2}}=1$. Given $\alpha>0$, we shall prove the following
$\Big|\{x:\mathcal{M}^{\mathbf{n}}(f_{1},f_{2})(x)>\alpha\}\Big|\leq C|\mathbf{n}|^{2}\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\alpha^{-\frac{2}{3}}.$
We apply the Calderón-Zygmund decomposition to $f_{1}$ at height $\max\left\{2^{|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\alpha^{\frac{2}{3}}$ to obtain the decomposition $f_{1}=g_{1}+h_{1}$ with a collection $\{Q_{\gamma}\}$ of disjoint cubes such that
$\|g_{1}\|_{L^{\infty}}\leq\max\left\{2^{|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\alpha^{\frac{2}{3}},~~h_{1}=\sum_{\gamma}h_{\gamma},~~supp(h_{\gamma})\subset Q_{\gamma},$
$\|h_{\gamma}\|_{L^{1}}\lesssim\max\left\{2^{|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\alpha^{\frac{2}{3}}|Q_{\gamma}|~~\text{ with}~~\int_{Q_{\gamma}}h_{\gamma}=0,~~\text{and}$
$\sum_{\gamma}\big|Q_{\gamma}\big|\leq\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\alpha^{-\frac{2}{3}}.$
@@
$|\{x\in\mathbb{R}^{d}:\mathcal{M}^{\mathbf{n}}(g_{1},f_{2})(x)>\frac{\alpha}{2}\}|\lesssim\frac{2^{-|\mathbf{n}|\frac{2d-1-2\beta}{2}}\|g_{1}\|_{L^{2}}\|f_{2}\|_{L^{2}}}{\alpha}.$
Since $\|g_{1}\|_{2}\lesssim\max\left\{2^{|\mathbf{n}|\frac{d-2-3\beta}{6}},2^{|\mathbf{n}|\frac{d-3-\beta}{6}}\right\}\alpha^{\frac{1}{3}}$, we obtain
$|\{x\in\mathbb{R}^{d}:\mathcal{M}^{\mathbf{n}}(g_{1},f_{2})(x)>\frac{\alpha}{2}\}|$ $\lesssim\max\left\{2^{-|\mathbf{n}|\frac{5d-1-3\beta}{6}},2^{-|\mathbf{n}|\frac{5d-5\beta}{6}}\right\}\alpha^{-\frac{2}{3}}$
$\lesssim\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\alpha^{-\frac{2}{3}}.$
@@
Then, $|\mathcal{F}|\leq\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\alpha^{-\frac{2}{3}}$. Thus, by Chebyshev’s inequality, it is enough to show that
$\int_{\mathbb{R}^{d}\setminus\mathcal{F}}|\mathcal{M}^{\mathbf{n}}(h_{1},f_{2})(x)|^{\frac{2}{3}}~dx\lesssim|\mathbf{n}|^{2}\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}.$ (6.2)
@@
$\int\limits_{\mathbb{R}^{d}\setminus\mathcal{F}}\Big|\mathcal{M}^{\mathbf{n}}_{k}(h^{i}_{1},f_{2})(x)\Big|^{\frac{2}{3}}dx\lesssim n_{2}\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\min\left\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\right\}\|h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|f_{2}\|^{\frac{2}{3}}_{L^{2}}.$
@@
$\int\limits_{\mathbb{R}^{d}\setminus\mathcal{F}}\Big|\mathcal{M}^{\mathbf{n}}_{k}(h^{i}_{1},f_{2})(x)\Big|^{\frac{2}{3}}dx$
$\lesssim n_{2}\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\min\left\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\right\}\|h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{\frac{2}{3}}_{L^{2}}.$
@@
$\lesssim\sum_{k,i\in{\mathbb{Z}}}n_{2}\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\min\Big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\Big\}\|h^{i}_{1}\|^{\frac{2}{3}}_{L^{1}}\|\mathbf{R}_{k+n_{2}}f_{2}\|^{\frac{2}{3}}_{L^{2}}$
$\lesssim n_{2}\min\left\{2^{-|\mathbf{n}|\frac{d-2-3\beta}{3}},2^{-|\mathbf{n}|\frac{d-3-\beta}{3}}\right\}\sum_{k\in{\mathbb{Z}}}\bigg(\sum_{i\in{\mathbb{Z}}}\min\Big\{1,2^{\frac{2(i-k)}{3}},2^{\frac{2(k+n_{1}-i)}{3}}\Big\}\bigg)^{\frac{1}{3}}$